Q.3 : Let X, Y be two random numbers with joint distribution function 

f(x,y) = Kx foryxy+1,0≤y≤2

= 0 otherwise. 

(a) Compute K

(b) Compute the m.d.f. and the expectation of X


1
Expert's answer
2021-07-13T12:58:28-0400

1= "\intop" "\intop" +ϖ f(x,y)dxdy ="\int" 02("\int" yy+1kxdx)dy= k/2"\intop" 02 ((y+1)2 -y)dy= k"\intop" 02 (y+ 0.5)dy

k/2(y+0.5)2 from 0 to 2

= k/2 (2.52 -0.52) =3k

Therefore k =1/3

Mean

E(x) ="\iint" xf(x,y)dxdy =1/9"\intop"02((y+1)3-y3)dy = 1/36(y+1)4 -y4) from 0 to 2

= (34 -24 -14 +0)/36

=16/9

The probability density function of x is fx(x) ="\intop" f(x,y)dy.

If x ∈/ (0,3)

fx(x)= 0 otherwise

fx(x) = "\intop" (0, x-1)(x,2) x/3.